Credit Risk Modeling: Default Probabilties

Jaime Frade December 19, 2008

Contents

1 Introduction 1 2 Preliminaries 2 3 Historical Models 3 4 Methodology 4 5 Data Description 5 6 Results and Predictions 6 7 Validation Techniques 7 8 Conclusion 8

List of Figures

ii

List of Tables

iii

Acknowledgements

iv Abstract

Chapter 1

Introduction

Credit Risk Methodology

In recent months more emphasis has shifted to the modeling and evaluation of credit risk. There are several forces behind this trend. First, credit markets have grown steadily and the credit derivatives market (CDS) has grown exponentially. Second, credit risk is still a developing field. The understanding and the methodologies in managing and measuring other market risks, including interest rate risk, currency risk etc., have matured and are now assessed based on widely accepted principles. Credit markets are implying significant levels of of default over the next twelve months, but rating agencies, such as Moody’s or Standard and Poors, downgrades and realized defaults have yet to follow suit. This is apparent in the latest subprime bubble bursts and the more recent defaults such as Lehman Brothers that have surprised the economy.

Chapter 2

Preliminaries

Chapter 3

Historical Models

Chapter 4

Methodology

Chapter 5

Data Description

Chapter 6

Results and Predictions

Chapter 7

Validation Techniques Chapter 8 Conclusion

8

Bibliography

9